OpenStudy (anonymous):

-2(x + 5) = -2x - 10 is an example of which property?

OpenStudy (agent47):

you are distributing the -2 in

OpenStudy (skullpatrol):

$$\Huge -2(x + 5) = -2x - 10$$ $$\Huge a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

Note that -2 is distributed as a multiplier of each term of the sum x +5.

OpenStudy (skullpatrol):